OpenStudy (anonymous):

Let g(x)= -x^(-2) + 6x^14 + pi^7. Find g'(0.5).

myininaya (myininaya):

\[g'(x)=-(-2)x^{-2-1}+6(14)x^{14-1}+0\] i used the power rule and the constant rule recall: \[(x^n)'=nx^{n-1} ; (c)'=0\]

OpenStudy (anonymous):

Do I need to find the deriv as the first step

myininaya (myininaya):

yes it ask for g'

OpenStudy (anonymous):

what about the qoutient rule for this problem?

myininaya (myininaya):

so plug in .5 into g'( )

myininaya (myininaya):

we don't have a quotient

OpenStudy (anonymous):

oh neveremind

OpenStudy (anonymous):

I got g'(x)+ 2x^-3 +84x^13

OpenStudy (anonymous):

what's next?

OpenStudy (anonymous):

im stuck

myininaya (myininaya):

plug in .5 and please don't post this twice if you need help fast then post in chat

myininaya (myininaya):

that is suppose to say g'(x)= not g'(x)+

OpenStudy (anonymous):

plug 0.5 for (x)?

OpenStudy (anonymous):

so

OpenStudy (anonymous):

g(x)= 2(0.5)^-3+84(0.5)^13

myininaya (myininaya):

that is g'(0.5)

OpenStudy (anonymous):

ok I got 1

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no 1+42^13